OpenStudy (anonymous):

Prime factorization of the lowest common multiple of 120 and 252

OpenStudy (unklerhaukus):

\[120=12\times10=2\times6\times2\times5=2\times2\times3\times2\times5=2^3\times3\times5\]

OpenStudy (unklerhaukus):

\[252=2\times126=2^2\times63=2^2\times9\times7=2^2\times3^2\times7\]